OpenStudy (anonymous):

Marine Biologists capture, tag, and release 137 Florida manatees into a protected region. Several months later, of 90 captured manatees, 6 have tags. Estimate the population of Florida manatees in the protected region. Show work

OpenStudy (kropot72):

The percentage of tagged manatees in the sample is: \[\frac{6}{90}\] Therefore 137 tagged manatees represent approximately 6/90 of the population. The estimated total population is:\[137\times \frac{90}{6}=?\]

OpenStudy (anonymous):

2055?

OpenStudy (kropot72):

Sorry I should have said "The fraction of tagged manatees........."

OpenStudy (kropot72):

Yes! The estimated population is 2055 :)